OpenStudy (anonymous):

Which of the following is the solution of log ( base x + 3) 0.001 = –3

OpenStudy (anonymous):

5,7, 13, 0r 15 are choices. please show me steps

OpenStudy (mathstudent55):

Use the definition of log. log(base b) x = y means b^y = x Here you have base = x + 3, so (x + 3)^-3 = 0.001, but 0.001 = 10^-3, so (x + 3)^-3 = 10^-3 If two powwers are equal and their exponents are equal, then the bases must be equal. x + 3 = 10 x = 7